OpenStudy (fibonaccichick666):

Prove that \(z=tan(\frac{\pi}{10})\) is a root of \(5z^5-5iz^4-10z^3-10iz^2+z-i=0\). I simplified it to \((5z^4-10z^2+1)(z-i)=0\), but I'm just not sure from here.

OpenStudy (fibonaccichick666):

So, if we plug it into the quadratic: \[z^2=\frac{10\left(\begin{matrix}+ \\ -\end{matrix}\right) \sqrt{80}}{10}\]

OpenStudy (fibonaccichick666):

So we get the 4 roots from \[z=\left(\begin{matrix}+ \\ -\end{matrix}\right) \sqrt{\frac{10\left(\begin{matrix}+ \\ -\end{matrix}\right) \sqrt{80}}{10}}\]

OpenStudy (fibonaccichick666):

@satellite73 , any ideas?

OpenStudy (anonymous):

hmm i am still at your step, solving the 4th degree equation i wonder if there is some half or double angle identity to use any clue? as in what the topic for the class is?

OpenStudy (anonymous):

oh i have an idea not saying it is right write it in polar form and see if it checks

OpenStudy (anonymous):

did you try working with \[5(z^2-1)^2-4=0\]?

OpenStudy (anonymous):

not sure that helps \[(z^2-1)^2=\frac{4}{5}\] \[z^2-1=\pm\frac{2}{\sqrt5}\]

OpenStudy (anonymous):

at least that proves that \(z^2\) is real, not sure where to go from there

myininaya (myininaya):

How do you get that factorization?

myininaya (myininaya):

I think there is something off about it unless I multiplied and divided wrong.

myininaya (myininaya):

I'm talking about what fibs did

OpenStudy (anonymous):

\[5z^4-10z^2+1=5(z^24-2z^2)+1=5(z^4-2z^2+1)+1-5=5(z^2-1)^2-4\]

OpenStudy (anonymous):

ok typo there, but you get the idea not sure how it help, but is seems to nice to be a coincidence

myininaya (myininaya):

was it suppose to be \[5z^5-5iz^4-10z^3+10iz^2+z-i=0\]?

OpenStudy (anonymous):

ooh are you saying the initial factorization is wrong? i didn't check ti

myininaya (myininaya):

yep i said fibs maybe i should have spelled out Fibonacci

myininaya (myininaya):

I think he meant that one part to be + instead - with it being - there is actually no real solutions (i cheated and used wolfram)

OpenStudy (anonymous):

oh i see lets wolfram

OpenStudy (anonymous):

btw, i assume that even though fibonacci was a guy, fibonacci chick is probably a woman

myininaya (myininaya):

oops i don't check last names my bad

OpenStudy (anonymous):

anyway i am stuck

myininaya (myininaya):

so I guess we need to somehow prove \[\tan(\pi/10)=\sqrt{1-\frac{2}{\sqrt{5}}}\]

myininaya (myininaya):

if we can prove this then maybe we can work backwards

OpenStudy (anonymous):

it is right, according to the wolf

OpenStudy (anonymous):

lol want to work through this? http://math.stackexchange.com/questions/196067/how-to-calculate-the-exact-value-of-tan-frac-pi10

OpenStudy (fibonaccichick666):

so sorry guys, i was afk

OpenStudy (anonymous):

i bet there is a snappier way, but i sure don't know it probably write the solution as \(z^2=1-\frac{2}{\sqrt5}\) and then make some argument (pun intended) about what \(z^2\) has to be in polar form

OpenStudy (fibonaccichick666):

topic is complex variables

OpenStudy (anonymous):

lol, got that

OpenStudy (anonymous):

also there must be a typo in the question before the factorization but no matter we got to \[(z^2-1)^2=\frac{4}{5}\] making \[z^2=1\pm\frac{2}{\sqrt5}\] which does in fact make \(z=\tan(\frac{\pi}{10})\)

OpenStudy (fibonaccichick666):

Sorry again, I was thinking about applying Cauchy Integral Theorem somehow

OpenStudy (fibonaccichick666):

the issue is pi,

OpenStudy (anonymous):

@Zarkon was here a second ago, he teaches this stuff

OpenStudy (fibonaccichick666):

but so, let me check real quick, the factorization may be wrong, but the initial problem is right

OpenStudy (anonymous):

i think the initial is wrong and the factorization is right, because the solution is right to the factored one

myininaya (myininaya):

That is what I was going to say. Took the words out of my mouth.

OpenStudy (fibonaccichick666):

nope, the initial is right, I just checked it against the problem

myininaya (myininaya):

That would lead to all imaginary solutions then

myininaya (myininaya):

therefore tan(pi/10) could not possibly be a root

OpenStudy (fibonaccichick666):

it's weird I know That's what is partially messing with me

myininaya (myininaya):

\[5z^5-5iz^4-10z^3+10iz^2+z-i=0 \] This is what I'm thinking what was meant.

OpenStudy (anonymous):

not possible wolf has 5 complex solutions to the original http://www.wolframalpha.com/input/?i=+5z^5%E2%88%925iz^4%E2%88%9210z^3%E2%88%9210iz^2%2Bz%E2%88%92i%3D0. but exactly the right solution to the factored one http://www.wolframalpha.com/input/?i=tan%28pi%2F10%29

OpenStudy (fibonaccichick666):

hmm if we redistribute the z-i which do we get, then let's go with that and say my teacher wrote the problem wrong

myininaya (myininaya):

When you multiply it you get the equation I wrote above.

OpenStudy (fibonaccichick666):

Let's say it's wrong and you are right(because most likely you are)

myininaya (myininaya):

There is just a one single difference between the questions there is a + in front of the 10iz^2

OpenStudy (fibonaccichick666):

I agree, what if I wrote the factorization wrong and it's a plus?

myininaya (myininaya):

well that factorization you wrote is correct for the + version

OpenStudy (fibonaccichick666):

then I don't get the minus i

OpenStudy (fibonaccichick666):

ok so i'm convinced,

OpenStudy (fibonaccichick666):

I was incorrect, thank you for that... it explains so much

myininaya (myininaya):

But I'm not sure if I can actually be help to this problem. I would have to do some reading.

OpenStudy (anonymous):

so far i am at \[\cos(2\theta)=1\pm\frac{2}{\sqrt5}\]

OpenStudy (fibonaccichick666):

actually, I think @satellite73 's factorization might do the trick, if I can prove that \(tan(\pi/10)=\sqrt{1(+/-)\frac2 {\sqrt5}}\)

OpenStudy (anonymous):

you can i have provided a link to the proof

OpenStudy (anonymous):

http://math.stackexchange.com/questions/196067/how-to-calculate-the-exact-value-of-tan-frac-pi10 but i bet there is another way, using complex numbers as this method only uses real numbers

OpenStudy (fibonaccichick666):

then, obviously it's a root right?

OpenStudy (fibonaccichick666):

well, i was thinking about using CIT and then having a be tan(pi/10)

OpenStudy (fibonaccichick666):

making it =0

OpenStudy (fibonaccichick666):

then using squeeze to equal zero

OpenStudy (fibonaccichick666):

but i just don't know how to apply it

OpenStudy (anonymous):

i would go with just using numbers but i could be wrong without a similar example worked out, i am not sure how to proceed

OpenStudy (fibonaccichick666):

i like yours better

OpenStudy (fibonaccichick666):

a lot better

OpenStudy (fibonaccichick666):

do you mind best answering @myininaya for me @satellite73 , she deserves a medal too

OpenStudy (anonymous):

i bet @Zarkon has a better method

OpenStudy (fibonaccichick666):

we'll see when he gets back can you check mine in a min?

OpenStudy (fibonaccichick666):

http://math.stackexchange.com/questions/7695/how-to-prove-cos-frac2-pi-5-frac-1-sqrt54 I want to use the complex way, but I don't understand the series

OpenStudy (zarkon):

I don't see a slick way of doing it. What you did is fine

OpenStudy (fibonaccichick666):

How did they write z=1+z+z^2+z^3+z^4=0?

OpenStudy (fibonaccichick666):

Thank you @Zarkon

OpenStudy (fibonaccichick666):

oops \(tan\theta=\frac{z-z^{-1}}{z+z^{-1}}\)

OpenStudy (fibonaccichick666):

\(z=e^\frac{i\pi}{10}\)

OpenStudy (fibonaccichick666):

@satellite73 , do you know how they did the series?

OpenStudy (fibonaccichick666):

ok i got the expansion, but why can i make the expansion?